THE B̃^{2}Σ_{u}^{+}−X̃^{2}Π_{g} SYSTEM OF SiNSi: INVESTIGATION OF ROVIBRONIC STRUCTURE AND RENNER-TELLER INTERACTION BY LASER INDUCED FLUORESCENCE SPECTROSCOPY.

\underline{Z. J. JAKUBEK}, M. ZACHWIEJA^{a}, S. G. NAKHATE^{b}, and B. SIMARD, Steacie Institute for Molecular Sciences, National Research Council, 100 Sussex Drive, Ottawa, Ontario K1A 0R6, Canada.

The free SiNSi radical has been investigated in a molecular free jet by laser induced fluorescence excitation and dispersed fluorescence spectroscopy. The radicals have been generated by laser ablation of Si in the presence of NH_{3} diluted (5%) in He. Excitation spectra of the B̃^{2}Σ_{u}^{+} − X̃^{2}Π_{g} system have been recorded in the 32600 − 36400 cm^{−1} range. For most of the excitation bands, dispersed fluorescence spectra have been obtained. Very large Renner-Teller interaction has been observed in both electronic states. The excitation and dispersed fluorescence spectra have been vibrationally and rotationally analyzed. Rotational, vibrational, and Renner-Teller parameters will be presented.

aOn leave from: Institute of Physics, University of Rzeszow, 35-310 Rzeszow, Poland

bOn leave from: Spectroscopy Division, Bhabha Atomic Research Centre, Mumbai, Maharastra 400 085 India